OpenStudy (howard-wolowitz):

Am I correct here with 2/5

OpenStudy (jhannybean):

First you'd have to begin by finding r.

OpenStudy (howard-wolowitz):

I see

OpenStudy (jhannybean):

You're given x, given y, and \[r=\sqrt{x^2+y^2}\]

OpenStudy (jhannybean):

By graphing your points, you'll see that it forms a right triangle. Thats when you use your trig ratios to find whatever you're looking for. In this case it's \(\tan(\theta)\).

OpenStudy (jhannybean):

So first of all, what is r = ?

OpenStudy (howard-wolowitz):

29

OpenStudy (jhannybean):

rather, \(r=\sqrt{29}\).

OpenStudy (jhannybean):

Good.

OpenStudy (jhannybean):

Now do you know SOHCAHTOA?

OpenStudy (howard-wolowitz):

yes

OpenStudy (jhannybean):

You're looking for \(\tan(\theta)\). Which part of SOHCAHTOA does that follow?

OpenStudy (howard-wolowitz):

H or T if I remeber

OpenStudy (jhannybean):

Hmm.. not quite. SOH : \(\sin\theta = \dfrac{opp}{hyp}\) CAH: \(\cos\theta = \dfrac{adj}{hyp}\) TOA: \(\tan\theta = \dfrac{opp}{adj}\)

OpenStudy (howard-wolowitz):

I said T

OpenStudy (jhannybean):

I wasn't quite sure what you were referring to with 'H'

OpenStudy (howard-wolowitz):

yes my mistake

OpenStudy (jhannybean):

But yes, \(\tan\theta = \dfrac{opp}{adj}\).

OpenStudy (jhannybean):

Now we draw out our triangle and fill in the values.

OpenStudy (howard-wolowitz):

person your a good teacher

OpenStudy (jhannybean):